OpenStudy (anonymous):

2. Write the equation of the line which passes through (2, 1) and is perpendicular to x = –2.

OpenStudy (anonymous):

The slope of a line x = -2 is not defined or infinity because line is perpendicular to x-axis therefore the slope of required line will be be zero The equation of the required line is y = 0(x) +C This passes through the point (2,1) y= 0(2) +1 y = 1